# ProxSparse: Regularized Learning of Semi-Structured Sparsity Masks for Pretrained LLMs

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional performance in natural language processing tasks, yet their massive size makes serving them inefficient and costly. Semi-structured pruning has emerged as an effective method for model acceleration, but existing approaches are suboptimal because they focus on local, layer-wise optimizations using heuristic rules, failing to leverage global feedback. We present ProxSparse, a learning-based framework for mask selection enabled by regularized optimization. ProxSparse transforms the rigid, non-differentiable mask selection process into a smoother optimization procedure, allowing gradual mask exploration with flexibility. ProxSparse does not involve additional weight updates once the mask is determined. Our extensive evaluations on 7 widely used models show that ProxSparse consistently outperforms previously proposed semi-structured mask selection methods with significant improvement, demonstrating the effectiveness of our learned approach towards semi-structured pruning.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated strong performance across a wide range of natural language processing (NLP) tasks (Brown, 2020; Achiam et al., 2023; Wei et al., 2022). However, deploying and serving these LLMs is not cost-efficient due to their massive size (Frantar & Alistarh, 2023; Yuan et al., 2024), often containing billions of parameters. To address the high computational demands and improve accessibility, various techniques have been proposed to make LLMs more efficient, such as model compression (Han et al., 2015; Frantar et al., 2022). By reducing memory footprint and accelerating computation, model compression significantly improves the feasibility and cost-effectiveness of deploying LLMs at scale (Liu et al., 2024b; Frantar et al., 2022; Lin et al., 2024).
@@
Network pruning (LeCun et al., 1989) is commonly used to reduce model size by removing unimportant parameters, lowering computational cost and improving efficiency (Bai et al., 2024). Among various pruning patterns, semi-structured pruning (Mishra et al., 2021), or block-wise N:M sparsification, has emerged as a practical and effective approach for LLM compression (Sun et al., 2023; Fang et al., 2024). In this approach, only N non-zero elements are retained out of M consecutive elements within each parameter block. This semi-structured sparsity strikes a balance between model accuracy and hardware efficiency, and is well-supported by many hardware accelerators (Mishra et al., 2021), enabling efficient serving for LLMs.
@@
Despite its advantages, finding an effective semi-structured mask for LLMs remains challenging. Pruning must follow a per-block structural restriction, making efforts on other patterns hard to adopt. Additionally, extensive retraining after pruning is impractical due to LLMs’ massive parameter size. (Ma et al., 2023). Recent advances like Wanda (Sun et al., 2023) and SparseGPT (Frantar & Alistarh, 2023) have improved semi-structured pruning using minimal resources with only hundreds of calibration samples, but still struggle to maintain optimal performance after pruning. We identify two main challenges in finding effective semi-structured masks: 1. The heuristic rules used for mask selection cannot fully take advantage of the calibration dataset during pruning. Methods like SparseGPT and Wanda rely on the Hessian matrix and importance scores to select elements to prune, but these lightweight criteria fail to effectively leverage or learn from the calibration data. 2. Both methods focus on solving a “local” optimization problem associated with individual layer, without considering the broader, end-to-end optimization across the entire model. In those methods, pruning is based on localized information within each layer, without considering the connections across layers. Thus they cannot benefit from the global feedback, limiting the overall effectiveness of the pruning method.
@@
We advocate a learning based solution for semi-structured mask selection that incorporates global feedback. We propose ProxSparse, which learns to discover semi-structured masks through an end-to-end optimization process, rather than solely relying on local, heuristic-based decisions. ProxSparse enables a finetuning-like procedure that learns the mask through only hundreds of calibration datasets with low resource utilization. The core of ProxSparse is the mask selection regularizer applied during learning, which transforms the rigid, non-differentiable mask-selection problem into a gradual search process. ProxSparse progressively enforces semi-structured sparsity and frozen weight constraints during training, gradually shrinking unimportant weights to be pruned. ProxSparse does not involve additional weight updates after deterimining the mask. One challenge in regularized learning is the efficiency of the solver, as a slow solver makes end-to-end learning on LLMs impractical. To address this, we developed a fast solver using iterative soft-thresholding, enabling efficient end-to-end learning at LLM scale.
@@
To comprehensively evaluate our method, we conducted extensive experiments on 7 widely used high-performance open-source models from four model families including Mistral (Jiang et al., 2023), Qwen (Yang et al., 2024), OpenLlama (Geng & Liu, 2023) and an anonymous model family (elaborated in Sec 4.1). The benchmarks cover language modeling and seven widely used natural language reasoning tasks. The results show that our regularized learning significantly outperforms baselines consistently accross all evaluated models, producing more effective pruning masks. Our contributions are summarized as follows:
@@ 3 Methodology
We introduce ProxSparse, a learning-based pruning method guided by a mask selection regularizer that generates high-quality semi-structured masks for efficient LLM serving. ProxSparse enables mask exploration in a global perspective by leveraging the gradient-based method, taking into account cross-layer connections with end-to-end feedback, rather than relying on localized, heuristic-based approaches for abrupt pruning. In this work, we focus specifically on 2:4 sparsity, and we discuss the extension to other sparsity patterns in Appendix F.
@@ 3.1 Relaxation and Structure-inducing regularization
$\text{Reg}_{2:4}$ decomposes into every 4-parameter block, where we apply the following regularizer (Anonymous, 2024) to enforce the sparse pattern.
@@ 3.3 Efficient Proximal Operator
This is a non-convex optimization problem. Anonymous (2024) showed that it can be solved with three convex subproblems.
Theorem 4 ((Anonymous, 2024)).
@@
This result suggests that we can simply enumerate the three candidate solutions and return the one with the smallest objective value. Anonymous (2024) thus proposed to solve for the “3-sparse” and “dense” solutions using interior point method (IPM) with a log-determinant barrier function, leading to the EnumIPM algorithm, which optimally solves (7). However, EnumIPM incurs high computational cost (Table 1). A faster heuristic, EnumPGD, was introduced to replace IPM with projected gradient descent without imposing semidefinite constraints. While EnumPGD improves efficiency, it sacrifices provably guarantees.
@@ 4.1 Models, tasks and baselines
We evaluated ProxSparse on four most advanced and widely used open-source LLM families: Mistral (Jiang et al., 2023), Qwen (Yang et al., 2024), OpenLlama (Geng & Liu, 2023) and an anonymous model family (Anon.Model). The specific models used in our experiments include Mistral-v0.1-7b, Mistral-v0.3-7b, Qwen2.5-14b, OpenLlama-7b-v2 and three models from the Anon.Model family. The three anonymous models used in our experiments belong to the same model family. We designate the three distinct LLMs in our study as Anon.Model-1, Anon.model-2, and Anon.model-3, arranged in increasing size and each with billions of parameters. The larger size of Anon.model-3 makes it comparatively easier to prune than the smaller models.
We assess the performance of pruned models from different pruning mechanisms on both zero-shot tasks and language modeling. For calibration, we followed Wanda (Sun et al., 2023) and SparseGPT (Frantar & Alistarh, 2023) to utilize the C4 (Raffel et al., 2020) dataset for calibration. Zero-shot performance was evaluated with the EleutherAI LM-Eval-Harness (Gao et al., 2024) on seven widely used tasks (Liu et al., 2024a), while Wikitext (Merity et al., 2016) perplexity (PPL) was used as the language modeling metric, consistent with previous evaluation protocol (Sun et al., 2023; Frantar & Alistarh, 2023). The experiments use 400 data samples for calibration unless specified, with consistent counts across baselines for fair comparison. We discuss MaskLLM and present ablation studies on mask effectiveness with regards to calibration sample size in Section 4.3.4. For hyperparameters and configurations, we detail them in Appendix B. Our experiments were done on Nvidia A100 GPUs.
@@ 4.2 End to end performance evaluation
Table 2: Experimental results on Wikitext perplexity (PPL) and 7 commonly used zero-shot natural language reasoning tasks comparing ProxSparse to 3 other baselines on 7 widely used LLMs (Anon.Model-2 results are deferred to Table 9). Bold indicates the best pruning performance, while italic represents the original unpruned performance. SparseGPT updates weights to minimize reconstruction error, while the other methods keep retained weights frozen. ProxSparse consistently yields better results compared to all other baselines.
@@
We first evaluate language modeling. ProxSparse surpasses magnitude pruning and outperforms Wanda, the SOTA mechanism without weight updates at the same scale. More specifically, ProxSparse achieves a PPL of 9.91 vs. Wanda’s 13.81 on OpenLlama-7b-v2 with 28% improvement. Similarly, ProxSparse achieves a PPL of 8.51 on Anon.Model-1, compared to Wanda’s 11.42, reflecting a 35% improvement. In the Anon.model-2 experiments shown in Table 9, ProxSparse reduces PPL from Wanda’s 20.91 to 13.63. Even when compared to SparseGPT, which updates the weights to minimize error, ProxSparse still outperforms it by up to an 18% margin, as demonstrated in the Anon.model-3 experiments. In summary, across different models, ProxSparse consistently achieves better PPL with a significant gap compared to other baselines.
@@ 4.3.2 The sparsity pattern enforcer
To better understand this phenomenon, we analyze the regularizer’s impact in detail. We show the evolution of 2:4 sparsity across different $\lambda_{1}$ values for OpenLlama-7b-v2 (Figures 1) and Anon.Model-1(Figure 4 in Appendix E) with consistent trend. We use OpenLlama-7b-v2 as the example. if $\lambda_{1}$ is too low, the model remains largely dense, as shown in Figures 4(a), (b) and (c). This suggests under-learning, where unimportant weights are not fully recognized by the end of learning, resulting in incomplete mask selection. Conversely, a high $\lambda_{1}$ value leads to early commitment to a specific mask. In Figure 4(d), the yellow line shows similarity to the ”early mask” obtained after just 10% of learning. The final mask retains $\sim$99.5similarity to the early one, indicating stalled optimization. In between, a balanced strength allows flexible mask exploration that avoids premature commitment, while also enabling more effective learning than excessively low values.
@@ 4.3.4 Performance evolution with varying numbers of calibration samples
Our method enables effective semi-structured mask selection with only hundreds of samples. Here we analyze performance based on the number of calibration samples with OpenLlama-7b-v2 and Anon.Model-1. We compare our results with MaskLLM, SparseGPT, and Wanda using 100, 200, and 400 samples. MaskLLM struggles with small sample sizes, making it a complementary method to ours in large-scale learning. We use the statistics reported in the MaskLLM paper (Fang et al., 2024). As shown in Table 6, MaskLLM performed worst on Anon.Model-1 with low sample sizes, likely due to its reliance on extensive training for effective masks. SparseGPT and Wanda showed minimal improvement with increased calibration samples, consistent with previous observations (Fang et al., 2024; Sun et al., 2023). ProxSparse achieved the best across these sample sizes, with slight performance gains as samples increased within our target range. This confirms the effectiveness of our method in learning towards an optimal mask for semi-structured sparsity.
@@
cells = c, hlines, vline2,5-6 = -, OpenLlama-7b-v2 & 100 200 400 Anon.Model-1 100 200 400
@@ 4.4 Improved efficiency during inference
Finally, we evaluate the efficiency-related metrics of the sparsified model produced by ProxSparse. We present wall-clock inference speedup and memory footprint improvements for the 2:4 semi-structured sparsified model induced by ProxSparse. Our experiments are conducted on Nvidia A100 GPUs. We utilize the Nvidia CUTLASS library as the underlying implementation for 2:4 semi-structured sparse operations.
@@ 5 Conclusion
Large language models (LLMs) excel in natural language processing tasks and other downstreaming tasks. However, LLMs suffer from high computational costs due to their enormous parameter sizes. Semi-structured sparsity can improve inference efficiency, though it remains challenging due to the structural constraints during pruning. We propose a learning-based method with regularized optimization, progressively explores optimal mask through end-to-end gradient feedback. Extensive evaluation shows that ProxSparse significantly outperforms previous methods, enabling better accuracy for LLM pruning, making model deployment more cost-effective.
@@ Appendix B Hyperparameters and configurations
Anon.Model-1 0.25 0 0.0001 Adamw 0.1
Anon.Model-2 0.5 0.5 0.0001 Adamw 0.1
Anon.Model-3 0.85 0 5.00E-05 Adamw 0.1
@@
Appendix D End-to-end evaluation results on Anon.Model-2
Table 9: Experimental results on Wikitext perplexity (PPL) and performance across 7 commonly used zero-shot natural language reasoning tasks comparing ProxSparse to 3 other baselines on Anon.Model-2. Bold indicates the best pruning performance, while italic represents the original unpruned performance. SparseGPT updates weights to minimize reconstruction error, while the other methods keep retained weights frozen. Similar to the results in Table 2, ProxSparse consistently yields better results compared to all other baselines.
@@
In this section, we further discuss the evaluation results from Table 2, focusing on Anon.Model-2. We present results on Wikitext perplexity (PPL) and performance across seven commonly used zero-shot natural language reasoning tasks, comparing ProxSparse to three baselines in Table 9. In the Anon.Model-2 experiments, ProxSparse significantly reduces perplexity from Wanda’s 20.91 to 13.63. The overall results, compared to Magnitude Pruning, Wanda, and SparseGPT, follow the same trends discussed in Section 4.2, with ProxSparse consistently outperforming all other baselines.
Appendix E Evolution of 2:4 sparsity across $\lambda_{1}$ on Anon.Model-1

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional performance in natural language processing tasks, yet their massive size makes serving them inefficient and costly. Semi-structured pruning has emerged as an effective method for model acceleration, but existing approaches are suboptimal because they focus on local, layer-wise optimizations using heuristic rules, failing to leverage global feedback. We present ProxSparse, a learning-based framework for mask selection enabled by regularized optimization. ProxSparse transforms the rigid, non-differentiable mask selection process into a smoother optimization procedure, allowing gradual mask exploration with flexibility. ProxSparse does not involve additional weight updates once the mask is determined. Our extensive evaluations on 7 widely used models show that ProxSparse consistently outperforms previously proposed semi-structured mask selection methods with significant improvement, demonstrating the effectiveness of our learned approach towards semi-structured pruning. Code available here.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated strong performance across a wide range of natural language processing (NLP) tasks (Achiam et al., 2023; Wei et al., 2022). However, deploying and serving LLMs is not cost-efficient due to their massive size with billions of parameters (Frantar & Alistarh, 2023; Sui et al., 2025). To address the high computational demands and improve accessibility, various techniques have been proposed to make LLMs more efficient, such as model compression (Han et al., 2015; Frantar et al., 2022). By reducing memory footprint and accelerating computation, model compression significantly improves the feasibility and cost-effectiveness of deploying LLMs at scale (Yuan et al., 2024; Lin et al., 2024; Tseng et al., 2025; Ozkara et al., 2025; Wei et al., 2025).
@@
Network pruning is commonly used to reduce model size and lower computation cost by removing unimportant parameters (Bai et al., 2024). Among various pruning patterns, semi-structured pruning (Mishra et al., 2021), or block-wise N:M sparsification, has emerged as a practical and effective approach for LLM compression (Sun et al., 2023; Fang et al., 2024). In this approach, only N non-zero elements are retained out of M consecutive elements within each parameter block. This semi-structured sparsity strikes a balance between model accuracy and hardware efficiency, and is well-supported by many hardware accelerators (Mishra et al., 2021), enabling efficient LLM serving.
@@
Despite its advantages, finding an effective semi-structured mask for LLMs remains challenging. Pruning must follow per-block structural restriction, making efforts on other patterns hard to adopt. Additionally, extensive retraining after pruning is impractical due to LLMs’ massive size (Ma et al., 2023; Chuang et al., 2024). Recent advances like Wanda (Sun et al., 2023) and SparseGPT (Frantar & Alistarh, 2023) improved semi-structured pruning using minimal resources with only hundreds of calibration samples, but still struggle to maintain optimal performance after pruning. We identify two main challenges in finding effective semi-structured masks: 1. The heuristic rules used for mask selection cannot fully take advantage of the calibration dataset during pruning. Methods like SparseGPT and Wanda rely on the Hessian matrix and importance scores to select elements to prune, but these lightweight criteria fail to effectively leverage or learn from the calibration data. 2. Both methods focus on solving a “local” optimization problem associated with individual layer, without considering the broader, end-to-end optimization across the entire model. In those methods, pruning is based on localized information within each layer, without considering the connections across layers. Thus they cannot benefit from the global feedback, limiting the effectiveness of the pruning method.
@@
We advocate a learning based solution for semi-structured mask selection that incorporates global feedback. We propose ProxSparse, which learns to discover semi-structured masks through an end-to-end optimization process, rather than solely relying on local, heuristic-based decisions. ProxSparse enables a finetuning-like procedure that learns the mask through only hundreds of calibration datasets with low resource utilization. The core of ProxSparse is the mask selection regularizer applied during learning, which transforms the rigid, non-differentiable mask-selection problem into a gradual search process. ProxSparse progressively enforces semi-structured sparsity and frozen weight constraints during training, and gradually shrinks unimportant weights to be pruned. ProxSparse does not involve additional weight updates after determining the mask. One challenge in regularized learning is the efficiency of the solver, as a slow solver makes end-to-end learning on LLMs impractical. To address this, we developed a fast solver using iterative soft-thresholding, enabling efficient end-to-end learning at LLM scale.
@@
To comprehensively evaluate our method, we conducted extensive experiments on 7 widely used high-performance open-source models from four model families including Mistral (Jiang et al., 2023), Qwen (Yang et al., 2024), OpenLlama (Geng & Liu, 2023) and Llama (Touvron et al., 2023) family. The benchmarks cover language modeling and seven widely used natural language reasoning tasks. The results show that our regularized learning significantly outperforms baselines consistently accross all evaluated models, producing more effective pruning masks. Our contributions are summarized as follows:
@@ 3 Methodology
We introduce ProxSparse, a learning-based pruning method guided by a mask selection regularizer that generates high-quality semi-structured masks for efficient LLM serving. ProxSparse enables mask exploration in a global perspective by leveraging the gradient-based method, taking into account cross-layer connections with end-to-end feedback, rather than relying on localized, heuristic-based approaches for abrupt pruning. In this work, we focus specifically on 2:4 sparsity, and we discuss the extension to other sparsity patterns in Appendix G.
@@ 3.1 Relaxation and Structure-inducing regularization
$\text{Reg}_{2:4}$ decomposes into every 4-parameter block, where we apply the following regularizer (Kübler et al., 2025) to enforce the sparse pattern.
@@ 3.3 Efficient Proximal Operator
This is a non-convex optimization problem. Kübler et al. (2025) showed that it can be solved with three convex subproblems.
Theorem 4 ((Kübler et al., 2025)).
@@
This result suggests that we can simply enumerate the three candidate solutions and return the one with the smallest objective value. Kübler et al. (2025) thus proposed to solve for the “3-sparse” and “dense” solutions using interior point method (IPM) with a log-determinant barrier function, leading to the EnumIPM algorithm, which optimally solves (7). However, EnumIPM incurs high computational cost (Table 1). A faster heuristic, EnumPGD, was introduced to replace IPM with projected gradient descent without imposing semidefinite constraints. While EnumPGD improves efficiency, it sacrifices provably guarantees.
@@ 4.1 Models, tasks and baselines
We evaluated ProxSparse on four most advanced and widely used open-source LLM families: Mistral (Jiang et al., 2023), Qwen (Yang et al., 2024), OpenLlama (Geng & Liu, 2023) and an Llama (Touvron et al., 2023) family. The specific models used in our experiments include Mistral-v0.1-7b, Mistral-v0.3-7b, Qwen2.5-14b, OpenLlama-7b-v2, Llama-2-7b, Llama-2-13b and Llama-3.1-8b.
We assess the performance of pruned models from different pruning mechanisms on both zero-shot tasks and language modeling. For calibration, we followed Wanda (Sun et al., 2023) and SparseGPT (Frantar & Alistarh, 2023) to utilize the C4 (Raffel et al., 2020) dataset for calibration. Zero-shot performance was evaluated with the EleutherAI LM-Eval-Harness (Gao et al., 2024) on seven widely used tasks (Liu et al., 2024), while Wikitext (Merity et al., 2016) perplexity (PPL) was used as the language modeling metric, consistent with previous evaluation protocol (Sun et al., 2023; Frantar & Alistarh, 2023). The experiments use 400 data samples for calibration unless specified, with consistent counts across baselines for fair comparison. We discuss MaskLLM and present ablation studies on mask effectiveness with regards to calibration sample size in Section 4.3.4. Comparisons on additional pruning mechanisms (ADMMPrune (Boža, 2024), OWL (Yin et al., 2023) and AlphaPrune (Lu et al., 2024)) are further detailed in Appendix E. For hyperparameters and configurations, we detail them in Appendix B. Our experiments were done on Nvidia A100 GPUs.
@@ 4.2 End to end performance evaluation
Table 2: Experimental results on Wikitext perplexity (PPL) and 7 commonly used zero-shot natural language reasoning tasks comparing ProxSparse to 3 other baselines on 7 widely used LLMs (Llama-3.1-8b results are deferred to Table 9). Bold indicates the best pruning performance, while italic represents the original unpruned performance. SparseGPT updates weights to minimize reconstruction error, while the other methods keep retained weights frozen. ProxSparse consistently yields better results compared to all other baselines.
@@
We first evaluate language modeling. ProxSparse surpasses magnitude pruning and outperforms Wanda, the SOTA mechanism without weight updates at the same scale. More specifically, ProxSparse achieves a PPL of 9.91 vs. Wanda’s 13.81 on OpenLlama-7b-v2 with 28% improvement. Similarly, ProxSparse achieves a PPL of 8.51 on Llama-2-7b, compared to Wanda’s 11.42, reflecting a 35% improvement. In the Llama-3.1-8b experiments shown in Table 9, ProxSparse reduces PPL from Wanda’s 20.91 to 13.63. Compare to the Llama-2-7b model, Llama-3.1-8b have more information encoded in the model weights as much larger training corpus was used during pretraining. This significant performance improvement highlights the potential of ProxSparse’s effectiveness in handling dense model pruning mask selection. Even when compared to SparseGPT, which updates the weights to minimize error, ProxSparse still outperforms it by up to an 18% margin, as demonstrated in the Llama-2-7b experiments. In summary, across different models, ProxSparse consistently achieves better PPL with a significant gap compared to other baselines.
@@ 4.3 Deep dive into the regularizing mechanism
Figure 1: Evolution of sparsity ratio on Llama-2-7b based on the degree of regularization. (a) Evolution of the 2:4 sparsity ratio over learning progress, where an insufficient regularization degree leads to under-learning. (b) With a larger $\lambda_{1}$ parameters shrink more quickly towards 2:4 sparsity, resulting in early commitment to a suboptimal mask. (c) Comparison of the 2:4 sparse block ratios at early (0.1 epochs) and final stages of learning. (d) Mask similarity between the final mask and the early mask obtained after 10% epochs of learning. An excessively large $\lambda_{1}$ results in premature mask commitment, causing mask selection to stagnate and hindering optimal mask discovery.
@@ 4.3.2 The sparsity pattern enforcer
To better understand this phenomenon, we analyze the regularizer’s impact in detail. We show the evolution of 2:4 sparsity across different $\lambda_{1}$ values for Llama-2-7b (Figures 1) and OpenLlama-7b-v2 (Figure 4 in Appendix F) with consistent trend. We use Llama-2-7b as the example. if $\lambda_{1}$ is too low, the model remains largely dense, as shown in Figures 4(a), (b) and (c). This suggests under-learning, where unimportant weights are not fully recognized by the end of learning, resulting in incomplete mask selection. Conversely, a high $\lambda_{1}$ value leads to early commitment to a specific mask. In Figure 4(d), the yellow line shows similarity to the ”early mask” obtained after just 10% of learning. The final mask retains $\sim$99.5similarity to the early one, indicating stalled optimization. In between, a balanced strength allows flexible mask exploration that avoids premature commitment, while also enabling more effective learning than excessively low values.
@@ 4.3.4 Performance evolution with varying numbers of calibration samples
Our method enables effective semi-structured mask selection with only hundreds of samples. Here we analyze performance based on the number of calibration samples with OpenLlama-7b-v2 and Llama-2-7b. We compare our results with MaskLLM, SparseGPT, and Wanda using 100, 200, and 400 samples. MaskLLM struggles with small sample sizes, making it a complementary method to ours in large-scale learning. We use the statistics reported in the MaskLLM paper (Fang et al., 2024). As shown in Table 6, MaskLLM performed worst on Llama-2-7b with low sample sizes, likely due to its reliance on extensive training for effective masks. SparseGPT and Wanda showed minimal improvement with increased calibration samples, consistent with previous observations (Fang et al., 2024; Sun et al., 2023). ProxSparse achieved the best across these sample sizes, with slight performance gains as samples increased within our target range. This confirms the effectiveness of our method in learning towards an optimal mask for semi-structured sparsity.
@@
cells = c, hlines, vline2,5-6 = -, OpenLlama-7b-v2 & 100 200 400 Llama-2-7b 100 200 400
@@ 4.4 Improved efficiency during inference
Finally, we evaluate the efficiency metrics of the sparsified model produced by ProxSparse. We present wall-clock inference speedup and memory footprint improvements for the 2:4 semi-structured sparsified model induced by ProxSparse. Our experiments are conducted on Nvidia A100 GPUs. We utilize the Nvidia CUTLASS library as the underlying implementation for 2:4 semi-structured sparse operations.
@@ 5 Conclusion
LLMs excel in natural language processing tasks and downstreaming tasks. However, they suffer from high computational costs due to the enormous parameter sizes. Semi-structured sparsity can improve inference efficiency, though it remains challenging due to the structural constraints during pruning. We propose a learning-based method with regularized optimization, progressively explores optimal mask through end-to-end gradient feedback. Extensive evaluation shows that ProxSparse significantly outperforms previous methods, enabling better accuracy for LLM pruning, making model deployment more cost-effective.
Impact statement
This paper presents work with the goal is to advance the field of machine learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix B Hyperparameters and configurations
Llama-2-7b 0.25 0 0.0001 Adamw 0.1
Llama-2-13b 0.5 0.25 0.0001 Adamw 0.1
Llama-3.1-8b 0.85 0 5.00E-05 Adamw 0.1
@@
Appendix D End-to-end evaluation results on Llama-3.1-8b
Table 9: Experimental results on Wikitext perplexity (PPL) and performance across 7 commonly used zero-shot natural language reasoning tasks comparing ProxSparse to 3 other baselines on Llama-3.1-8b. Bold indicates the best pruning performance, while italic represents the original unpruned performance. SparseGPT updates weights to minimize reconstruction error, while the other methods keep retained weights frozen. Similar to the results in Table 2, ProxSparse consistently yields better results compared to all other baselines.
@@
In this section, we further discuss the evaluation results from Table 2, focusing on Llama-3.1-8b. We present results on Wikitext perplexity (PPL) and performance across seven commonly used zero-shot natural language reasoning tasks, comparing ProxSparse to three baselines in Table 9. In the Llama-3.1-8b experiments, ProxSparse significantly reduces perplexity from Wanda’s 20.91 to 13.63. The overall results, compared to Magnitude Pruning, Wanda, and SparseGPT, follow the same trends discussed in Section 4.2, with ProxSparse consistently outperforming all other baselines.
Appendix E Comparison with Additional Pruning Baselines (ADMMPrune, OWL, and AlphaPrune)
Table 10: Experimental results on Wikitext perplexity (PPL) and 7 commonly used zero-shot natural language reasoning tasks comparing ProxSparse to 3 other baselines Llama-2-7b and Mistral-v0.3-7b model. Bold indicates the best pruning performance, while italic represents the original unpruned performance. AdmmPrune updates weights to minimize reconstruction error, while OWL and AlphaPrune uses dynamic sparse ratios across layers. ProxSparse consistently yields better results compared to all other baselines.
{tblr}
cells = c, vline2-3,11 = -, hline1,2,7,12 = -, & Weight Update Wikitext PPL ARC-C ARC-E SIQA HellaSwag OBQA PIQA TruthfulQA AVG
OWL ✗ 13.03 0.275 0.594 0.406 0.417 0.188 0.688 0.320 0.413
AlphaPrune ✗ 13.58 0.265 0.529 0.398 0.407 0.190 0.668 0.335 0.399
ADMMPrune ✓ 9.06 0.340 0.680 0.416 0.471 0.240 0.739 0.299 0.455
ProxSparse ✗ 8.68 0.362 0.697 0.429 0.525 0.242 0.751 0.321 0.475
Llama-2-7b - 5.12 0.433 0.763 0.461 0.571 0.314 0.781 0.321 0.521
OWL ✗ 13.17 0.287 0.591 0.407 0.420 0.228 0.695 0.339 0.425
AlphaPrune ✗ 13.01 0.293 0.607 0.406 0.411 0.238 0.690 0.317 0.424
ADMMPrune ✓ 9.67 0.328 0.653 0.413 0.440 0.248 0.714 0.302 0.442
ProxSparse ✗ 8.51 0.331 0.656 0.407 0.478 0.242 0.716 0.328 0.452
In this section, we compare ProxSparse with three additional pruning baselines: ADMMPrune (Boža, 2024), OWL (Yin et al., 2023), and AlphaPrune (Lu et al., 2024). The experiments are done on Mistral-v0.3-7b and Llama-2-7b model. ADMMPrune introduces a fast and effective weight update algorithm for layerwise pruning based using the Alternating Direction Method of Multipliers (ADMM). As shown in table 10, ProxSparse outperforms ADMMPrune in both models, achieving lower PPL (8.51 vs. 9.67) and higher acc (47.6% vs. 45.5%), highlighting its effectiveness. We attribute the superority of ProxSparse to its end-to-end optimization process, which goes beyond solely relying on local layer-wised information.
OWL and AlphaPrune aim to determine layer-specfic ratio to protect important layers. Here we argue that they are not the best-suited mechanism in semi-structured pruning, as the sparse operator supported by hardware typically requires all blocks to strictly adhere the pattern, making applying varying ratios hard. Nevertheless, we conduct experiments on AlphaPrune and OWL for comparison. We follow mixed sparsity proposed in OWL and AlphaPrune with Wanda, that layers can have varying ratios, while the overall ratio remains 2:4. We see ProxSparse outperforms OWL and Alphaprune on Llama-2-7b and Mistral-v0.3-7b on PPL and accuracy, showing the strength of our end-to-end optimization. Further, as pruning patterns become more fine-grained (e.g., 2:4), varying layer-wise pruning ratios become less effective as critical weights might still be removed within each block. This was reported in (Yin et al., 2023; Lu et al., 2024), where 4:8 pruning performed just similarly to uniform pruning in Wanda. This highlights the benefits of ProxSparse in identifying fine-grained semi-structured masks.
Appendix F Evolution of 2:4 sparsity across $\lambda_{1}$ on OpenLlama-7b-v2